A solid sphere is rolling on a rough horizontal surface with linear speed v . It collides with a vertical wall elastically. Coefficient of friction on each surface is 0.5 .Find speed of sphere when it again starts rolling without slipping.


1
Expert's answer
2021-01-13T11:36:29-0500

"F_fr=I\alpha"


"I=\frac{2}{5}mr^2"


"\alpha=\frac{5F_f}{2mr}"


After collision we have the equation


"\omega_f=-\omega+\alpha t\to\omega_f=-\omega+\frac{5F_f}{2mr} t"


"v_f=-v+r\alpha t\to v_f=-v+\frac{5F_f}{2m} t"


"v_f=-v+\frac{5}{2}(v-v_f)\to v_f=\frac{3}{7}v" . Answer